During the last second of its free fall a body covers half of the total distance traveled . Initial velocity is zero . Find the height from which the body falls.

Grade:11

1 Answers

Naveen Kumar
6 years ago
Let total time taken by the body is t seconds.
 
Therefore distance travelled in t seconds :
S = ut + 1/2 at2
S= 0 + 1/2(10t2)
S= 5n2.
 
Distance traveled in last second i.e. in tth second.
S= u + a/2(2t - 1)
S= 0 + 10/2 (2t -1)
S= 10t - 5.
 
Also given :
St = S/2
10t - 5 = 5n/2
5n2 - 20t + 10 = 0
From here t= 2 seconds.
Putting value of t for S.
S= 5t2
S= 5 (2)²
S= 20 m.